in a mixture of concrete there are three lb of cement mix for each pound of gravel .if the mixture contains a total of 76 lb of these two ingredients how many pounds of gravel are there?
Found 3 solutions by josgarithmetic, macston, lwsshak3:
Answer by josgarithmetic(39618)

Fraction as cement: 3/4
Fraction as gravel: 1/4

Cut the 76 into four equal parts.

Answer by macston(5194)

.
G=amount of gravel; C=amount of cement=3G
.
G+C=76 lbs
G+3G=76 lbs
4G=76 lbs
G=19 lbs
.
ANSWER: There are 19 lbs of gravel in 76 lbs of mix.
.

Answer by lwsshak3(11628)

in a mixture of concrete there are three lb of cement mix for each pound of gravel .if the mixture contains a total of 76 lb of these two ingredients how many pounds of gravel are there?
let x=amt of gravel in mixture
76-x=amt of cement in mixture
(76-x)/x=3
3x=76-x
4x=76
x=19
76-x=57
amt of gravel in mixture=19 lbs